Present value. Prestigious University is offering a new admission and tuition payment plan for all alumni. On the birth of a child, parents can guarantee admission to Prestigious if they pay the first years tuition. The university will pay an annual rate of return of 6% on the deposited tuit on, and a full re und will be available fthe child chooses another university. The tuition is expected to be $22,000 a year at Prestigious 19 years from now. What would parents pay today if they just gave birth to a new baby and the child will attend college in 19 years? How much is the required payment to secure admission for their child if the interest rate falls to 4%? What would parents pay today if they just gave birth to a new baby and the child will attend college in 19 years? (Round to the nearest cent.)
Answer #1

a) FV of tuition fees after 19 years = 22,000
PV of tuitions today = 22000/(1+6%)19 = 7271.29

The parents will pay today = 7271.29

b) at rate = 4%
PV of tuitions today = 22000/(1+4%)19 = 10422.13

The parents will pay today = 10422.13

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well